OpenStudy (anonymous):

how many different ways can you choose two vowels out of the box, if there are 5 vowels in the box and others dont matter

OpenStudy (ikram002p):

\(c(n,k)=\left(\begin{matrix}n \\ k\end{matrix}\right)=\dfrac{n!}{p!(n-k)!}\) \(c(5,3)=\left(\begin{matrix}5 \\ 3\end{matrix}\right)=\dfrac{5!}{3!(5-3)!}\)=5×4×3=60

OpenStudy (ikram002p):

is it the right answer ?

OpenStudy (anonymous):

YES THANK U

OpenStudy (anonymous):

NO ITS NOT

OpenStudy (ikram002p):

mmm :O why !

OpenStudy (anonymous):

ITS 60, 10, 20,120

OpenStudy (anonymous):

ITS MULTIPLE CHOICE

OpenStudy (ikram002p):

ok so 60 is wrong ?

OpenStudy (anonymous):

MY BAD NOT 60 ITS 6

OpenStudy (ikram002p):

ok i guess the answer would be ,mmm 10

OpenStudy (anonymous):

OK LET ME CHECK THANKS